The circles center is at (2,2)and contains (6,5). What is the length of the diameter
Answer by stanbon(75887)

The circles center is at (2,2)and contains (6,5). What is the length of the diameter
-------
radius = sqrt[(5-2)^2+(6-2)^2] = sqrt[9+16]= sqrt(25) = 5
---
diameter = 2*radius = 10
===========================
Cheers,
Stan H.